OpenStudy (anonymous):

How do I find the angle between vectors u: cos(pi/3) + sin(pi/3) and v: cos(3pi/12) + sin(3pi/12)

OpenStudy (nikvist):

Do you think \[\vec{u}=\cos\frac{\pi}{3}\vec{i}+\sin\frac{\pi}{3}\vec{j}\]\[\vec{v}=\cos\frac{\pi}{4}\vec{i}+\sin\frac{\pi}{4}\vec{j}\]